# RmGPT: Rotating Machinery Generative Pretrained Model

### v1

@@ 3.3 Training Process
The training process of the RmGPT model involves two main stages: pretraining and prompt learning. Each stage plays a crucial role in ensuring that the model can generalize well to various diagnosis and prognosis tasks.
3.3.1 Next Signal Token Prediction Pretraining
@@
We propose a pretraining strategy specifically designed for PHM applications, inspired by techniques from the natural language processing domain. The objective is to predict the next signal token in a sequence, which prevents the model from simply repeating patterns and forces it to learn the underlying dynamics of the system. Accurately predicting the next token indicates that the model has learned the intrinsic signal dynamics, which are crucial for representing the system health state. Formally, given a sequence of signal tokens $\{T_{s1},T_{s2},\ldots,T_{sn}\}$, the model is trained to predict the next token $T_{si+1}$ based on the preceding tokens $\{T_{s1},T_{s2},\ldots,T_{si}\}$:
$\mathcal{L}_{\text{pretrain}}=-\sum_{i=1}^{n}\log P(T_{si+1}\mid T_{s1},T_{s2},\ldots,T_{si};\theta_{e})$ (11)
where $\theta_{e}$ represents the pretraining parameters. This strategy effectively captures temporal dependencies and patterns within the signal data, providing a solid foundation for downstream tasks.
3.3.2 Efficient Prompt Learning
We also propose an efficient prompt learning approach to adapt the pretrained model to specific tasks with minimal additional training. This method introduces task-specific prompts into the model, guiding its adaptation process. The prompt tokens $T_{p}\in\mathbb{R}^{l_{p}\times M_{i}\times d}$ are added to the input sequence, effectively steering the model’s attention towards task-relevant features:
@@ 3.3.2 Efficient Prompt Learning
$\mathcal{L}_{\text{prompt}}=\sum_{j=1}^{m}\mathcal{L}_{\text{task}}(T_{sj}+T_{p};\theta_{p})$ (12)
where $\theta_{p}$ represents the parameters specific to prompt learningand $\mathcal{L}_{\text{task}}$ is the loss function for the downstream task.
The advantages of this approach are twofold: (1) By fine-tuning only a few learnable embeddings, the model can efficiently adapt to various tasks and diagnosis equipment, significantly broadening its application scenarios. (2) This method enhances the model generalization performance by making minimal adjustments to the pretrained parameters obtained from vast amounts of data.
The relationship between pretraining and prompt learning is integral to the model training process. pretraining provides the model with a robust initialization, capturing general patterns in the signal data. Prompt learning allows the model to quickly adapt to new tasks with minimal data, leveraging task-specific information encoded in the prompts. This comprehensive training process ensures that the RmGPT model remains flexible and effective across a wide range of diagnosis and prognosis tasks, adapting seamlessly to different datasets and conditions.
4 EXPERIMENTS
@@
In this section, we conduct a comprehensive evaluation of the RmGPT model across various diagnosis and prognosis tasks using diverse datasets. We begin by describing the datasets employed in our experiments, highlighting their diversity and relevance to real-world applications. Following this, we present the main results, comparing the performance of RmGPT to SOTA algorithms[20, 21, 22, 23, 24, 25, 26, 27]. Additionally, we assess the few-shot learning capabilities of RmGPT, demonstrating its effectiveness in scenarios with limited training data. Subsequently, we conduct an efficiency analysis of RmGPT during the training and inference processes. Finally, we analyze the token space representations learned by RmGPT, illustrating its ability to transform noisy input signals into a clear and distinct semantic space.
4.1 Dataset description
@@
In this study, we employ a variety of datasets to evaluate the performance of the RmGPT model on both diagnosis and prognosis tasks. These datasets, summarized in Table 1, originate from different types of equipment and exhibit a wide range of characteristics, including signal channels, fault categories, sampling frequencies and dataset sizes.
Table 1: Overview of datasets used for diagnosis and prognosis tasks
Tasks Object Datasets Classes Channels Total Points Sample Frequency
Diagnosis Bearing CWRU[28] 16 1, 2, 3 9,011,145 12 kHz
Bearing SLIET[29] 13 3 13,631,475 70 kHz
Gear QPZZ-II[30] 5 9 1,064,960 5120 Hz
Gear SMU[31] 3 1 300,000 10 kHz
Prognosis Bearing XJTU[32] 4 2 204,046,336 25.6 kHz
@@
4.2 Main results
In our experiments, we train the RmGPT model using a set of carefully selected hyperparameters to ensure optimal performance across various diagnosis and prognosis tasks. The experiments are conducted on a computing server equipped with 8 NVIDIA RTX 3090 GPUs.
To ensure consistency across diverse datasets, we uniformly downsampled all data to approximately 5kHz and standardized the input signal windows to 2048 time steps. This preprocessing step was essential to harmonize the datasets, allowing the model to learn underlying patterns without being influenced by variations in signal amplitude or sampling rates.
The hyperparameters employed for training the RmGPT model are outlined in Table 2. In the Tokenizer settings, both stride length and patch length were set to 256. The model architecture was designed with 4 layers and 512 hidden units, amounting to a total of 68.50M parameters, ensuring sufficient representational capacity. The prompt and fault token lengths were set to 10 and 1, respectively. During the optimization phase, the model was trained with a batch size of 256 and a learning rate of $3.00\times 10^{-7}$. The training regimen comprised 20 epochs of pretraining, conducted as self-supervised learning across all datasets without label information, followed by 3 epochs of fine-tuning and 5 epochs of prompt learning.
Table 2: Hyperparameters Used for Training RmGPT Model
@@ 4.2 Main results
Pretrain Epochs 20
Finetune Epochs 3
Prompt Epochs 5
Stride Length (S) 256
Patch Length (P) 256
Layers 4
Hidden Numbers (d) 512
@@
Model Size 68.50 M
We conducted a comprehensive evaluation of the RmGPT model across various diagnosis and prognosis tasks using multiple datasets. To ensure the fairness of all comparisons, the experimental setup maintained a consistent training-to-testing ratio of 8:2 across all datasets. For the RmGPT model, we conducted a single pretraining phase on all available training data, followed by supervised fine-tuning across all datasets simultaneously. This approach allows RmGPT to leverage the shared information across different datasets, enhancing its generalization capabilities. In contrast, the comparison SOTA models were individually fine-tuned on each dataset, optimizing their performance for specific tasks but lacking the unified approach inherent to RmGPT. This rigorous and consistent evaluation framework underscores the robustness and adaptability of RmGPT in diverse diagnostic and prognostic scenarios, further validating its superiority as a foundation model for rotating machinery. The results, summarized in Table 3, highlight one of the most significant contributions of our work: the ability to use a single model with the same parameters across all datasets while outperforming SOTA algorithms in diverse tasks.
In the diagnosis tasks, the RmGPT model consistently demonstrated superior performance, significantly surpassing all comparison models, particularly in challenging datasets like QPZZ and SLIET. The RmGPT-Prompt configuration achieved an impressive average accuracy of 99.21% across multiple datasets, with individual dataset accuracies ranging from 98.05% to 100%. The RmGPT-Finetune configuration further improved upon this, achieving an average accuracy of 99.75% and achieving perfect scores on several datasets.
In prognosis tasks, RmGPT also excelled, particularly in predicting the Remaining Useful Life (RUL) of rolling bearings, where it outperformed all comparison models. Specifically, the RmGPT-Prompt configuration achieved a Mean Absolute Error (MAE) of 0.150 and a Mean Squared Error (MSE) of 0.033, while the RmGPT-Finetune configuration further reduced these errors to 0.136 MAE and 0.030 MSE.
The use of a single set of PHM model parameters to effectively handle diverse datasets with significant variability demonstrates the feasibility of constructing a unified PHM model for fault diagnosis and prognosis. This unified approach paves the way for creating large-scale models that can generalize across various rotating machinery.
Table 3: RmGPT Uses the Same Parameter Model and Outperforms SOTA Algorithms on Different Tasks and Equipments
@@
Transformer[21] 99.39% 100.00% 91.02% 65.67% 89.02% 0.231 0.070
TICNN[25] 90.30% 90.00% 96.26% 66.67% 85.81% 0.181 0.043
CNN_LSTM[22] 86.73% 96.98% 93.85% 66.67% 86.05% 0.186 0.045
DP_MRTN[27] 100.00% 100.00% 98.33% 98.02% 99.09% 0.250 0.083
IMSFACNN[26] 98.62% 75.00% 98.11% 66.67% 84.60% 0.250 0.083
ResNet18[20] 99.71% 98.03% 97.66% 48.89% 86.07% 0.259 0.091
InversePINN[24] 99.48% 100.00% 96.29% 80.67% 94.11% 0.159 0.035
TARTDN[23] 91.07% 95.00% 97.94% 63.56% 86.89% 0.158 0.034
4.3 Few-Shot Learning Capabilities
@@
To further evaluate the adaptability and robustness of the RmGPT model, we conducted few-shot learning experiments on the CWRU dataset. In these experiments, the training samples were limited to a few examples per class. Specifically, 1-shot learning indicates that there is one sample per class for all 16 classes, resulting in a total of 16 training samples. Similarly, 4-shot learning means there are four samples per class, totaling 64 training samples, and so on. The training set was used without any data augmentation. The number of testing samples was kept constant at 13,734. Importantly, all methods, including RmGPT and the compared models, were evaluated without any prior exposure to additional samples from the CWRU dataset, ensuring that none of the models had seen the test data during training. Moreover, the training and testing sample sizes were kept consistent across all models to guarantee a fair comparison. This rigorous approach ensures that the observed performance differences genuinely reflect the models’ capabilities in few-shot learning scenarios.
@@
The results of the few-shot learning experiments are summarized in Table 4. RmGPT consistently outperforms other models, demonstrating its ability to achieve high accuracy with very few training samples. This superior performance can be attributed to the extensive pretraining of RmGPT on large datasets, which enables the model to learn generalizable fault signal patterns. The RmGPT model, particularly under the prompt learning configuration, shows remarkable performance even in extreme conditions. In the 1-shot scenario, RmGPT-Prompt achieves a diagnosis accuracy of 92.02% across 16 fault classes, significantly outperforming other models. In the 8-shot scenario, RmGPT-Prompt reaches an accuracy of 99.43%, further highlighting its robustness and effectiveness.These results demonstrate that RmGPT can efficiently leverage small amounts of data to achieve high-precision diagnostics. The model’s exceptional performance in few-shot scenarios makes it highly suitable for practical equipment maintenance situations where fault samples are often scarce. This validates the feasibility of applying a general-purpose model like RmGPT in real-world scenarios.
Table 4: Few-shot Diagnosis Performance on CWRU Dataset
@@
These results underscore the strength of RmGPT in practical scenarios, demonstrating its potential to be used as a universal model for various rotating machinery diagnostics and prognostics, even with limited fault data.
4.4 Efficiency Analysis
To evaluate the efficiency of the RmGPT model, we compared its performance under various configurations, focusing on inference time, FLOPS, backward time, and the number of training parameters. The results are summarized in Table 5.
The impact of the proposed patch-based tokenization and Time-Channel Attention mechanism on reducing computational complexity and improving performance is significant. The original configuration shows an inference time of 4.970 ms and FLOPS of 51.097 GFLOPS. Removing the patch-based tokenization increases the inference time to 337.773 ms and FLOPS to 3813.998 GFLOPS. Similarly, excluding the Time-Channel Attention mechanism results in an inference time of 9.281 ms and FLOPS of 102.214 GFLOPS. For training efficiency, the prompt learning approach is more efficient than finetuning. The backward time for prompt learning is 6.765 ms with 3.038 M training parameters, compared to 22.023 ms with 68.548 M parameters for finetuning.
Table 5: Efficiency Comparison of RmGPT

### v2

@@ III-C Training Process
The training process of the RmGPT model involves two main stages: pretraining and downstream task adaption. Each stage plays a crucial role in ensuring that the model can generalize well to various diagnosis and prognosis tasks.
III-C1 Next Signal Token Prediction Pretraining
@@
We propose a pretraining strategy specifically designed for PHM applications, inspired by techniques from the natural language processing domain. The objective is to predict the next signal token in a sequence, which prevents the model from simply repeating patterns and forces it to learn the underlying dynamics of the system. Accurately predicting the next token indicates that the model has learned the intrinsic signal dynamics, which are crucial for representing the system health state.
For the RmGPT model, the input is a patch-tokenized multivariate sequence $x_{p}\in\mathbb{R}^{l_{s}\times M\times P}$, where each of the $M$ channels has been segmented into $l_{s}$ temporal patches of length $P$. During pretraining, we mask the final patch $x_{p_{l_{s}}}$ of each channel and train the model to reconstruct it based on the preceding patches $\{x_{p_{1}},x_{p_{2}},\ldots,x_{p_{l_{s}-1}}\}$.
Formally, the pretraining objective is to minimize the reconstruction loss between the true final patch and the predicted output from the model:
$\mathcal{L}_{\text{pretrain}}=\left\|x_{p_{l_{s}}}-G(z_{l_{s}})\right\|^{2}$ (13)
where $z_{l_{s}}\in\mathbb{R}^{M\times d}$ denotes the predicted latent vector corresponding to the masked token position, and $G(\cdot)$ is a lightweight decoder that projects this latent vector back to the patch space $\mathbb{R}^{M\times P}$.
The latent vector $z_{l_{s}}$ is computed by applying the Transformer encoder $F(\cdot;\theta_{e})$ over the visible patch tokens:
$[z_{1},z_{2},\ldots,z_{l_{s}}]=F(x_{p_{1}},x_{p_{2}},\ldots,\texttt{[MASK]};\theta_{e})$ (14)
and extracting the output at the masked position $z_{l_{s}}$. In other words, the model is trained to infer the representation at the masked location from the surrounding context, and then reconstruct the raw signal patch through the decoder head.
This formulation encourages the model to learn informative temporal representations by reconstructing raw signal content from the masked token embedding, rather than directly generating the signal. In our implementation, we use a fixed-length context of $n=7$ patches to predict the 8th patch for each channel. This self-supervised design enables the model to capture localized dynamics and cross-channel dependencies in a semantically grounded and computationally efficient manner.
To adapt the pretrained RmGPT model to downstream diagnosis and prognosis tasks, we consider two complementary strategies: full finetuning and prompt learning.
III-C2 Prompt Learning
To efficiently adapt the pretrained model to specific diagnosis or prognosis tasks, we adopt a lightweight prompt tuning framework, which introduces a small number of learnable task-specific tokens into the model input while keeping the pretrained backbone parameters $\theta_{e}$ frozen. Specifically, we incorporate three types of tokens:
  • Prompt Tokens $T_{p}$: encode task- or equipment-specific information, guiding the model’s attention toward relevant signal features;
  • Task Tokens $T_{t}$: act as aggregation centers for multichannel signal representations, producing a global health descriptor;
  • Fault Tokens $T_{f}$: serve as semantic anchors for different fault types, enabling interpretable and similarity-based classification.
During training, only the parameters associated with these tokens—denoted as $\theta_{p}\subset\theta_{e}$—are updated, while the remaining pretrained parameters remain unchanged. This leads to a parameter-efficient adaptation strategy, requiring the tuning of less than 5% of the total model parameters.
The prompt-enhanced input is constructed as $(T_{s}+T_{p})$, and the training objective is to optimize only the prompt-related parameters $\theta_{p}$ based on task-specific supervision:
@@ III-C2 Prompt Learning
$\theta_{p}=\arg\min_{\theta_{p}\subset\theta_{e}}\mathbb{E}_{(\mathcal{X}_{i},\mathcal{Y}_{i})\sim D}\left[\mathcal{L}_{\text{task}}(\mathcal{Y}_{i},F(\mathcal{X}_{i}+T_{p};\theta_{p}))\right]$ (15)
This design enables flexible adaptation in scenarios with limited training data or constrained computational resources, while reducing the risk of overfitting. The task tokens $T_{t}$ play a central role in summarizing multichannel signal features into a global health state, which is compared against fault tokens $T_{f}$ to produce interpretable and discriminative predictions.
III-C3 Full-Parameter Finetuning
As an alternative to prompt learning, RmGPT also supports traditional finetuning, where all model parameters $\theta_{f}$ are updated using labeled data from the target task:
$\theta_{f}=\arg\min_{\theta_{f}}\mathbb{E}_{(\mathcal{X}_{i},\mathcal{Y}_{i})\sim D}\left[\mathcal{L}(\mathcal{Y}_{i},F(\mathcal{X}_{i};\theta_{f}))\right]$ (16)
Here, $\theta_{f}$ shares the same dimensionality and structure as the pretrained parameters $\theta_{e}$, and can be viewed as a continuously optimized version of $\theta_{e}$ tailored to the target task. In other words, full-parameter finetuning initializes from $\theta_{e}$ and updates the entire parameter set to better fit the new data distribution.
This approach enables deeper task-specific adaptation and potentially stronger performance when sufficient labeled data is available. However, it also introduces higher computational overhead and a greater risk of overfitting, particularly in low-resource settings.
III-C4 Adaptation Methods: Prompt vs. Finetuning
In contrast, prompt learning provides an efficient and lightweight alternative by updating only a small number of learnable task-specific tokens ($\theta_{p}$: $T_{p}$, $T_{t}$, $T_{f}$), while keeping the backbone frozen. This allows for fast adaptation across different tasks or domains, with lower risk of overfitting and reduced computational cost.
Table I summarizes the differences between these two strategies across key aspects, including parameter efficiency, expressiveness, and suitability for different scenarios.
TABLE I: Comparison of Prompt Learning and Full Finetuning in RmGPT
Aspect Prompt Learning Full Finetuning
Updated Parameters Only task-specific tokens ($\theta_{p}$: $T_{p}$, $T_{t}$, $T_{f}$) All model parameters ($\theta_{f}$)
Parameter Ratio Less than 5% 100%
Efficiency Fast, low computation Slow, high computation
Overfitting Risk Low High
Expressiveness Moderate (leveraging pretrained features) High (full model adaptation)
Use Cases Few-shot learning, cross-domain transfer Full adaptation with sufficient labeled data
Overall, the two strategies offer flexible trade-offs: prompt learning is more suitable for rapid deployment in resource-constrained environments, while full finetuning provides greater expressiveness and task-specific capacity when sufficient data and compute are available.
The overall training process first equips the model with general signal understanding through self-supervised pretraining, followed by lightweight or full adaptation to downstream tasks. Prompt learning enables efficient task transfer by updating only a small set of task-specific embeddings, while full finetuning allows complete adaptation when sufficient data is available. Together, these strategies make RmGPT a flexible and powerful foundation model for rotating machinery health management across varying equipment, signal modalities, and fault conditions.
IV EXPERIMENTS
@@
In this section, we conduct a comprehensive evaluation of the RmGPT model across various diagnosis and prognosis tasks using diverse datasets. We begin by describing the datasets employed in our experiments, highlighting their diversity and relevance to real-world applications. Following this, we present the main results, comparing the performance of RmGPT to SOTA algorithms[chang2023intelligent, sun2023liteformer, chen2021bearing, cui2024triplet, qin2024inverse, xu2021fault, chen2022dual]. Additionally, we assess the few-shot learning capabilities of RmGPT, demonstrating its effectiveness in scenarios with limited training data. Subsequently, we conduct an efficiency analysis of RmGPT during the training and inference processes. Finally, we analyze the token space representations learned by RmGPT, illustrating its ability to transform noisy input signals into a clear and distinct semantic space.
IV-A Dataset description
@@
In this study, we employ a variety of datasets to evaluate the performance of the RmGPT model on both diagnosis and prognosis tasks. These datasets, summarized in Table II, originate from different types of equipment and exhibit a wide range of characteristics, including signal channels, fault categories, sampling frequencies and dataset sizes.
TABLE II: Overview of datasets used for diagnosis and prognosis tasks
Tasks Object Datasets Classes Channels Points Frequency
Diagnosis Bearing CWRU[shi2024cross] 16 1, 2, 3 9,011,145 12 kHz
Bearing SLIET[SLIET] 13 3 13,631,475 70 kHz
Gear QPZZ-II[QPZZ] 5 9 1,064,960 5120 Hz
Gear SMU[SMU] 3 1 300,000 10 kHz
Prognosis Bearing XJTU[XJTU] 4 2 204,046,336 25.6 kHz
@@
IV-B Experimental Setup
All experiments are conducted on a computing server equipped with 8 NVIDIA RTX 3090 GPUs. To ensure consistency across diverse datasets, we uniformly downsample all signals to approximately 5kHz and standardize input windows to 2048 time steps. This preprocessing harmonizes variations in signal amplitude and sampling rates, enabling the model to focus on learning meaningful temporal patterns.
We adopt a unified training pipeline for all datasets. The training consists of two phases: a self-supervised pretraining stage and a supervised adaptation learning stage. In the first phase, the RmGPT model is pretrained for 20 epochs on the unlabeled training data aggregated from all datasets, enabling the model to learn generalizable temporal representations. In the second phase, we perform adaptation learning using labeled data, with two different adaptation strategies: prompt learning and full-parameter finetuning. The only difference between the two lies in the scope of trainable parameters—prompt learning updates only a small subset of task-specific tokens ($\theta_{p}\subset\theta_{e}$), whereas finetuning updates the entire parameter set ($\theta_{f}$) initialized from $\theta_{e}$. All other training configurations, including learning rate, optimizer, number of epochs, and batch size, are kept consistent across both adaptation strategies to ensure fair comparison.
The model architecture consists of 4 Transformer layers with 512 hidden units, totaling 68.50M parameters. This compact yet expressive configuration is selected to balance representation capacity and computational efficiency. The tokenizer uses a stride and patch length of 256, and the prompt and fault token lengths are set to 10 and 1, respectively.
Table III summarizes the hyperparameter settings used throughout all experiments.
TABLE III: Hyperparameters Used for Training RmGPT Model
@@ IV-B Experimental Setup
Pretraining Epochs 20
Finetuning Epochs 3
Prompt Learning Epochs 5
Tokenizer Stride (S) 256
Tokenizer Patch Length (P) 256
Transformer Layers 4
Hidden Size ($d$) 512
@@
Total Model Parameters 68.50M
IV-C Main results
We conducted a comprehensive evaluation of RmGPT across a wide range of diagnosis and prognosis tasks using multiple public rotating machinery datasets. To ensure fairness, all experiments adopt a consistent training-to-testing split ratio of 8:2. Following self-supervised pretraining on the aggregated unlabeled data, RmGPT is adapted to downstream tasks via two adaptation strategies: prompt learning and full-parameter finetuning. Except for the number of trainable parameters, all experimental settings are kept identical across these two adaptation modes. In contrast, all baseline models are trained individually on each dataset in a supervised manner, without benefiting from shared pretraining or a unified adaptation mechanism.
This rigorous and consistent evaluation framework highlights the robustness and versatility of RmGPT across heterogeneous PHM scenarios. A key advantage of our approach lies in the ability to use a single model configuration—with shared pretrained parameters and unified adaptation settings—across all datasets and tasks. As summarized in Table IV, RmGPT consistently outperforms existing state-of-the-art (SOTA) methods, demonstrating its effectiveness as a general-purpose foundation model for rotating machinery health management.
In fault diagnosis tasks, both RmGPT-Prompt and RmGPT-Finetune configurations show strong performance. The prompt-based version achieves an average accuracy of 99.21% across datasets, with individual accuracies ranging from 98.05% to 100%, while the finetuned version further improves to an average of 99.75%, achieving perfect scores on several challenging datasets such as QPZZ and SLIET.
In prognosis tasks, RmGPT also delivers superior results. For RUL prediction on rolling bearing datasets, the RmGPT-Prompt configuration achieves a Mean Absolute Error (MAE) of 0.150 and a Mean Squared Error (MSE) of 0.033. The RmGPT-Finetune configuration achieves even better performance, reducing the errors to 0.136 MAE and 0.030 MSE.
These results collectively demonstrate the feasibility and advantage of constructing a unified PHM model that can generalize across diverse equipment types, signal characteristics, and task objectives. The ability to deploy a single model with consistent architecture and training strategy across heterogeneous datasets marks an important step toward scalable, foundation-level modeling in industrial prognostics and health management.
TABLE IV: RmGPT Uses the Same Parameter Model and Outperforms SOTA Algorithms on Different Tasks and Equipments
@@ IV-C Main results
Transformer[sun2023liteformer] 99.39% 100.00% 91.02% 65.67% 89.02% 0.231 0.070
CNN_LSTM[chen2021bearing] 86.73% 96.98% 93.85% 66.67% 86.05% 0.186 0.045
DP_MRTN[chen2022dual] 100.00% 100.00% 98.33% 98.02% 99.09% 0.250 0.083
IMSFACNN[xu2021fault] 98.62% 75.00% 98.11% 66.67% 84.60% 0.250 0.083
ResNet18[chang2023intelligent] 99.71% 98.03% 97.66% 48.89% 86.07% 0.259 0.091
InversePINN[qin2024inverse] 99.48% 100.00% 96.29% 80.67% 94.11% 0.159 0.035
TARTDN[cui2024triplet] 91.07% 95.00% 97.94% 63.56% 86.89% 0.158 0.034
IV-D Few-Shot Learning Capabilities
@@
To further evaluate the adaptability and robustness of the RmGPT model, we conducted few-shot learning experiments on the CWRU dataset. In these experiments, the training samples were limited to a few examples per class. Specifically, 1-shot learning indicates that there is one sample per class for all 16 classes, resulting in a total of 16 training samples. Similarly, 4-shot learning means there are four samples per class, totaling 64 training samples, and so on. The training set was used without any data augmentation. The number of testing samples was kept constant at 13,734. Importantly, all methods, including RmGPT and the compared models, were evaluated without any prior exposure to additional samples from the CWRU dataset, ensuring that none of the models had seen the test data during training. Moreover, the training and testing sample sizes were kept consistent across all models to guarantee a fair comparison. This rigorous approach ensures that the observed performance differences genuinely reflect the models’ capabilities in few-shot learning scenarios. It is also important to note that while the CWRU dataset may be partially included in the unlabeled corpus used during the self-supervised pretraining stage, no label information was provided to the model at that time. The pretraining process relied solely on unlabeled signal sequences, without access to class identities or diagnostic labels. Therefore, the few-shot learning performance reflects the model’s ability to transfer general temporal representations learned from unlabeled data, rather than any memorization of labeled examples. This distinction ensures the validity of the few-shot learning evaluation and highlights the benefit of the foundation model’s generalization capability. The results of the few-shot learning experiments are summarized in Table V.
@@
RmGPT consistently outperforms other models, demonstrating its ability to achieve high accuracy with very few training samples. In the 1-shot scenario, RmGPT-Prompt achieves a diagnosis accuracy of 82.46% across 16 fault classes, significantly outperforming other models. In the 4-shot scenario, RmGPT-Prompt reaches an accuracy of 98.53%, further highlighting its robustness and effectiveness. These results demonstrate that RmGPT can efficiently leverage small amounts of data to achieve high-precision diagnostics. RmGPT’s exceptional performance in few-shot learning scenarios makes it highly suitable for practical equipment maintenance situations, where fault samples are often scarce.
These results underscore the strength of RmGPT in practical scenarios, demonstrating its potential to be used as a universal model for various rotating machinery diagnostics and prognostics, even with limited fault data. The model’s exceptional performance in few-shot learning scenarios can be attributed to several factors. First, RmGPT benefits from extensive pretraining on large datasets, which enables it to learn generalizable fault signal patterns. This pretraining helps the model extrapolate from minimal examples, making it highly effective in adapting to new fault categories with limited labeled samples. Second, the generative nature of RmGPT allows it to infer unseen fault patterns by leveraging its learned representations of signal semantics, a critical capability in few-shot settings where data is scarce. Lastly, RmGPT’s prompt learning mechanism facilitates rapid adaptation to new tasks by efficiently utilizing the few available training samples. This combination of pretraining, generative modeling, and efficient task adaptation makes RmGPT highly robust in few-shot learning scenarios, where limited data availability is often a challenge.
TABLE V: Few-shot Diagnosis Performance on CWRU Dataset
@@
IV-E Efficiency Analysis
To evaluate the efficiency of the RmGPT model, we compared its performance under various configurations, focusing on inference time, FLOPS, backward time, and the number of training parameters. The results are summarized in Table VI.
The impact of the proposed patch-based tokenization and channel-time Attention mechanism on reducing computational complexity and improving performance is significant. The original configuration shows an inference time of 4.970 ms and FLOPS of 51.097 GFLOPS. Removing the patch-based tokenization increases the inference time to 337.773 ms and FLOPS to 3813.998 GFLOPS. Similarly, excluding the channel-time Attention mechanism results in an inference time of 9.281 ms and FLOPS of 102.214 GFLOPS. For training efficiency, the prompt learning approach is more efficient than finetuning. The backward time for prompt learning is 6.765 ms with 3.038 M training parameters, compared to 22.023 ms with 68.548 M parameters for finetuning.
TABLE VI: Efficiency Comparison of RmGPT